OpenStudy (i_love_my_nieces):

On which of the following lines does (7,1) lie? y - 5x = 4 y = x - 10 y = -4x = 3 y = -x + 8

OpenStudy (i_love_my_nieces):

@undeadknight26

OpenStudy (i_love_my_nieces):

@austinL

OpenStudy (i_love_my_nieces):

@__Links__

undeadknight26 (undeadknight26):

I would say D because: y = -x + 8 1 = -7 + 8 It works out :D

OpenStudy (i_love_my_nieces):

Thanks bud!